OpenStudy (anonymous):

Anyone knows how to do this?? Separate 38 into two numbers such that three times the smaller is 16 less than twice the larger.

hartnn (hartnn):

yup.

hartnn (hartnn):

let 38 be separated into x and 38-x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do we set them equal to each other?

hartnn (hartnn):

three times the smaller is 16 less than twice the larger. so 3x+16 2(36-x) got this ?

OpenStudy (theeric):

Only if those two "numbers" are equal to each other! Otherwise, no!

hartnn (hartnn):

3x+16= 2(36-x)

OpenStudy (anonymous):

okay, then solve right?

hartnn (hartnn):

yup

hartnn (hartnn):

find x from there, u will get the smaller number

OpenStudy (anonymous):

then substitute for the right, correct?

OpenStudy (anonymous):

larger*

hartnn (hartnn):

yes, but first tell me what did u get 'x' value.

OpenStudy (anonymous):

11.2

OpenStudy (anonymous):

is that correct?

hartnn (hartnn):

yes, even i get 11.2, i wonder why is it not and integer, anyways, find larger using 36-x= ?

OpenStudy (anonymous):

using 36? or 38?

hartnn (hartnn):

sorry 38

OpenStudy (anonymous):

its okay, but I got 26.8 for large

hartnn (hartnn):

ohh.thats why we didn't get an integer, its 38, sorry my typing mistake

OpenStudy (anonymous):

oh

hartnn (hartnn):

solve this : 3x+16= 2(38-x)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i got 28 for large, 10 for small

hartnn (hartnn):

i got x=12, please verify the calculations.....

OpenStudy (anonymous):

oh yeah, your're right

OpenStudy (anonymous):

the larger is 26

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^